OpenStudy (sally):

I am having a brain fart, i need guidance.. 5^(3x-1)-2=23

OpenStudy (anonymous):

5^(3x-1)=25 5^(3x-1)=5^2 by comparing powers of 5. 3x-1=2 x=1

OpenStudy (sally):

Sam, thank you verymuch. you make it look so easy..